Exercises dA (1) Use Cauchys residue theorem to compute Jo 2+sin (2) Repeat the preceding exercise for 8 131. (3) Let a beplease 2 only, thanks
Answer #1

Date: Page 10 Problem2.. do I = 172+2 coso ? het z = elo __वर = lode limits RT Pizle iz 13+27 Z+z} 2 f... coso = e io te lotDate: 1 1 _ Page no: _ & idz Jit (32+z2+1) J. (2°+32+) IT Jizie) i (+2.612)7 Z+0.38) The residue theorem implies I= 2Ti residDate: 11 Page no: Vee z = -0.38 Eid (zto:28) Z (2-1), Idz itz +2.673)2 (2+0.3527 X=-0.38 f (x+2.618300) - (2) (2) (z+2.618)75-1 Date: 1 1 Page no:_ - (2.618-0.38) + 2 (0.38) į (2.618-0.38)3 2.238 +0.76 11.20 i _ 0.267 Tru, I - 2 TO. 0.267 1.677 ANSW